OpenStudy (anonymous):

8n - (-2)/7 = 6 I have no idea how to go about solving this.

OpenStudy (anonymous):

just think its -(-a) = a an identity. also use the cross multiplication for 7 in the denominator. and solve for n

OpenStudy (anonymous):

\[\frac{ 8n }{ 7 }=6 ; 8n= 6*7 ; n = \frac{ 42 }{8 }\]

OpenStudy (anonymous):

8n-(-2)/7=6 multiply by 7, 56n-(-2)=42 56n+2=42 56n=42-2=40 n=40/56=5/7

OpenStudy (anonymous):

Oh, I see, well thank you c: